# A Conditional Diffusion Model for Building Energy Modeling Workflows

### v1

@@ 2.1 Deep Learning in Energy Systems and Buildings application
Our proposed work differs fundamentally by developing a generative modeling framework instead of a predictive one. This is a more challenging and powerful approach, as our generative model learns the entire underlying conditional data distribution over multiple building characteristics, rather than a direct predictive mapping. We train a single, unified diffusion model with conditional generation capabilities that can be sampled to generate diverse plausible values for any number of missing building characteristics conditioned on the observed attributes. Our model is trained once on the comprehensive ResStock dataset, which represents the entire U.S. residential building stock. This enables it to generalize across diverse geographical locations, as opposed‘ to training a separate predictive neural network model for each district ([6, 5]).
2.2 Generative models for Tabular Dataset
Tabular datasets are ubiquitous in many domains, and developing generative models for tabular data has many important use cases [32, 19]. High-quality synthetic tabular generation could be used for data augmentation and is in demand for its privacy-preserving benefits. Generative models also provide capabilities for missing value imputation by framing it as a conditional generation task, which is the primary focus of our study. However, tabular data presents a challenge as it is often comprised of heterogeneous or mixed-type features, requiring the generative approach to jointly learn distributions over both discrete and continuous variables.
@@ 3.2 Conditional TabDDPM
The numerical features are transformed with the Gaussian quantile transformation [48], and the categorical features are ordinal encoded before being one-hot encoded. We use parameters $f_{\text{mask, num}}$ and $f_{\text{mask, cat}}$ to control for the maximum proportion of the numerical variables and categorical variables to be masked while training. During each training step, a random positive masking ratio between $(0,f_{\text{mask, num}})$ and $(0,f_{\text{mask, cat}})$ is selected, and the corresponding number of numerical and categorical variables (respectively) is masked or set unobserved for every row in the batch. Notably, while the same number of variables is masked from every row in a batch, the variables themselves are chosen at random. This is the dynamic masking strategy described above. To this end, for (1), we train a single model with $f_{\text{mask, num}}=0.5$ and $f_{\text{mask, cat}}=0.2$. For (2), we train two separate models with all numerical features observed and $f_{\text{mask, cat}}$ set to 0.05 and 0.4, respectively.
@@ 4 Evaluation
Dataset: Our study utilizes the ResStock [12] dataset (https://resstock.nrel.gov/). ResStock is a synthetic dataset statistically representative of the entire residential building stock in the U.S [26]. This dataset is generated by sampling from probability distributions derived from multiple data sources [12], resulting in an extensive collection of U.S residential buildings at PUMA-level resolution [6, 5]. A PUMA varies in area, consists of many geographically close buildings, and contains granular details on building characteristics in the area. We use the 2024.1 release version of the ResStock dataset, which comprises 2.2 million buildings. Each row of this tabular data corresponds to a unique building, and the columns include various building characteristics and energy consumption details. From this list of available features, we select a small subset to work with:
@@
In total, we have 35 features (4 numerical and 31 categorical). We always keep the latitude and longitude columns known or observed in our conditional generation setup. We use a small subset of withheld PUMAs [26] to create an out-of-distribution (OOD) test dataset for evaluating our generative model’s generalization capabilities. Overall, we split the dataset into train (1,405,212), validation (351,303), test (439,128), and OOD-test set (3,665).
@@ 4.1.1 Univariate analysis
Result summary: The resulting true and generated distributions are visualized as histograms, and we show plots for both numerical and categorical features in Figure 2. Quantitative distances are reported in Table 1, showing an average over every unique dependency set for each variable. The Wasserstein distance is normalized by the range of the true samples. Table 1 includes results for the mixed-imputation model. We also compute results for one of the categorical-only imputation models, trained using a masking ratio of $0.05$ (which is equivalent to masking up to $2$ out of $31$ categorical features during training). The performance of the categorical-only imputation model remains consistent with that of the mixed imputation model, with an average JS-Distance of $0.197$ after averaging over all categorical variables. For comparison, we include a baseline that imputes each categorical variable by randomly choosing a category from all its possible categories. The JS distance for this baseline is averaged across all variables and reported in the last row of Table 1, which shows the random baseline’s poor performance as compared to our model.
@@
We observe from Table 1 and Figure 2 that the model demonstrates strong conditional generation quality across many features, though the performance varies, with some variables performing better than others. According to Table 1, some variables, such as heating/cooling Setpoint has offset, heating fuel, and building geometry type, are particularly challenging for the model to learn. This may be due to complex underlying class structures or imbalanced classes for some of these variables [5]. Figure 2 shows that the model’s output is realistic and matches the true distribution closely, even with features having a higher number of categories, e.g., for refrigerator, vintage (built year range), and water heater efficiency; the generated distribution also captures the dominant modes well, like "Natural Gas Standard" for water heater efficiency. Notably, the model performs well even on features that lack strongly defined dependencies in the ResStock data, such as the clothes dryer usage level variable. The model is also effective in capturing the wide range of values present in the numerical distribution, e.g, for square footage.
@@
Analysis on OOD: Here, we validate our model’s generalization capability on the out-of-distribution (OOD) test dataset. Overall, we expect OOD conditional generation to be a challenging task, particularly for the variables that directly or indirectly depend on geospatial factors, as buildings in these PUMAs were not seen during training. The Figure 3 shows generated vs true distribution plots on a subset of the variables visualized in Figure 2. We see similar qualitative performance in Figure 3 for all these variables. It is important to note that for variables such as vintage or building square footage, which are dependent on PUMA (as seen in the dependencies obtained from ResStock), a comparison to the true distribution is not possible since there are no matching training samples with the same PUMA. However, this challenge highlights the advantage of our approach - learning a single generative model on the entire U.S. residential building stock. This enables our model to still generalize from broader patterns and impute building characteristics for geographical locations unseen during training, unlike studies that train a separate predictive neural network model for each district ([6, 5]).
@@ 4.3 Case Study
We demonstrate the application of our framework in a case study on a residential neighborhood in Baltimore, Maryland (see Figure 5). Our study integrates our conditional generative modeling pipeline with the URBANopt [7, 8] energy model following the setup in Kontar et al. [5]. Specifically, given a set of observed characteristics for the $77$ buildings in this neighborhood, we employ our Conditional TabDDPM model to impute all the unknown building characteristics. This synthetically completed data is then passed as an input to URBANopt, a physics-based energy model, which simulates load profiles in terms of energy or electricity consumption per building.
Figure 5: A residential community in Baltimore is used in our case study (based on [5]). There are $77$ single-family attached buildings. Building characteristics such as built year, number of stories, heating fuel, attic type, and building area (square footage) are obtained for these buildings from sources like OpenStreetMap and Zillow/Redfin.
We utilize building characteristics for the buildings in this neighborhood as curated in El Kontar et al.’s work [5], and we refer to this as the reference dataset. These include building attributes collected from open-source databases and characteristics produced by their neural network-based predictive model (discussed in 2.1). From this set of characteristics, we select the $35$ building features in the Conditional TabDDPM model setup. We keep all the numerical features as observed and mask varying levels of categorical features. We start by masking one categorical feature, then two features, and then progressively increase the number of masked features to $10$, to assess our model’s ability in imputing increasing levels of unknown characteristics conditioned on the rest of the observed attributes. After imputing the corresponding number of missing (masked) variables with our model, we evaluate the imputed features’ quality by comparing URBANopt’s resulting load profile using this generated dataset against the true load profile, which is simulated using the reference dataset from [5].
@@ 5 Discussion
The results presented in the paper successfully validate the use of a conditional diffusion framework in generating realistic building attributes and creating complete characteristics in building-level datasets. Our findings in Section 4 demonstrate the model’s ability to generate imputed features whose distributions closely match the true ResStock distribution for many building variables. However, a feature-level breakdown of the JS distance between the two distributions in Table 1 shows that a few variables are more challenging to learn than others. It is possible that some of these features, like building types and the number of occupants, which are critical conditioning variables for generating other variables [5], are not strongly determined by other observed features in return. Moreover, detailed analysis of some variables, including a few with the worst performance, such as heating/cooling setpoint has offset, heating fuel, and water heater fuel, reveals that our model learns an almost deterministic conditional mapping in these specific cases. The model consistently collapses to a single generated outcome when it is sampled conditioned on a given set of observed values, as seen in Figure 8. While these variables perform great at single-point prediction, scoring high in reconstruction accuracy, the probabilistic nature of the outputs is not captured well in these cases. Although we observe this behavior to be limited to only a few variables, it is possible that an improved diffusion model workflow could address this in future work.
Overall, our evaluation focused on comparing the generated and true probability distributions in the univariate and bivariate conditional generation setting. While we don’t extend this analysis to higher-dimensional multivariate cases, we demonstrate our model’s multivariate conditional generation capabilities by simultaneously imputing up to $10$ building attributes, i.e, nearly a third of the total variables, in a real-world case study evaluation. Furthermore, our validation experiments with the OOD dataset show the model’s strong generalization capability, which is an important advantage of our global generative modeling framework. Finally, we showcase an end-to-end application where our model generates a completed building-level dataset needed as an input to the URBANopt energy model. Our case study on a small residential community analyzes the simulated energy load profiles for these buildings, which closely match a curated (reference) baseline. This study underscores our model’s potential in generating physically plausible and reliable datasets for practical use cases, showing its utility in enhancing building energy modeling workflows. Future work can involve expanding case studies to more complex and larger communities.
@@ 6 Conclusion
Acknowledgments
This work was authored by the National Renewable Energy Laboratory (NREL), operated by Alliance for Sustainable Energy, LLC, for the U.S. Department of Energy (DOE) under Contract No. DE-AC36- 08GO28308. This work was supported by the Laboratory Directed Research and Development (LDRD) Program at NREL. The views expressed in the article do not necessarily represent the views of the DOE or the U.S. Government. The U.S. Government retains and the publisher, by accepting the article for publication, acknowledges that the U.S. Government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for U.S. Government purposes. The research was performed using computational resources sponsored by the Department of Energy’s Office of Energy Efficiency and Renewable Energy and located at the National Renewable Energy Laboratory.

### v2

# Conditional Distribution Estimation of Building Characteristics with Diffusion Models for Urban Energy Modeling

@@ 2.1 Deep Learning in Energy Systems and Buildings application
Our proposed work differs fundamentally by developing a generative modeling framework instead of a deterministic one. This is a more challenging and powerful approach, as our generative model learns the entire underlying conditional data distribution over multiple building characteristics, rather than a direct point mapping. We train a single, unified diffusion model with conditional generation capabilities that can be sampled to generate diverse plausible values for any number of missing building characteristics conditioned on the observed attributes. Our model is trained once on the comprehensive ResStock dataset, which represents the entire U.S. residential building stock. This enables it to be applied across diverse geographical locations, as opposed to training a separate neural network model for each district ([13, 11]).
2.2 Deep Learning for Tabular Data Imputation
Tabular datasets are ubiquitous in many domains. Classical statistical methods like MICE [46] provide a standard baseline for tabular data imputation by iteratively estimating missing values, but they scale poorly to large datasets and struggle to capture complex, non-linear feature interactions [38]. Many deep learning approaches for tabular data overcome these limitations, and treat imputation as a deterministic task, mapping available inputs to a single point estimate for the missing value [5]. Recent models utilizing masked autoencoding frameworks [9], contrastive learning [28], and implicit neural representations [34] have improved prediction accuracy on incomplete datasets. Concurrently, modern architectures like TabM [18] have refined how neural networks process tabular structures compared to traditional tree-based models (e.g. gradient-boosted decision tress [7]) which have historically dominated this domain [40]. However, these methods learn a point mapping to an expected value rather than capturing the underlying distribution of the missing data, highlighting the need for deep generative models that learn the joint distribution over features, as discussed below.
2.2.1 Generative models for Tabular Data
Developing generative models for tabular data has many important use cases [39, 27]. High-quality synthetic tabular generation could be used for data augmentation and is in demand for its privacy-preserving benefits. Generative models also provide capabilities for missing value imputation by framing it as a conditional generation task [52, 53], which is the primary focus of our study. However, tabular data presents a challenge as it is often comprised of heterogeneous or mixed-type features, requiring the generative approach to jointly learn distributions over both discrete and continuous variables.
@@ 3.2 Conditional TabDDPM
The numerical features are transformed with the Gaussian quantile transformation [35], and the categorical features are ordinal encoded before being one-hot encoded. We use parameters $f_{\text{mask, num}}$ and $f_{\text{mask, cat}}$ to control for the maximum proportion of the numerical variables and categorical variables to be masked while training. During each training step, a random positive masking ratio between $(0,f_{\text{mask, num}})$ and $(0,f_{\text{mask, cat}})$ is selected, and the corresponding number of numerical and categorical variables (respectively) is masked or set unobserved for every row in the batch. Notably, while the same number of variables is masked from every row in a batch, the variables themselves are chosen at random. This is the dynamic masking strategy described above. To this end, for (1), we train a single model with $f_{\text{mask, num}}=0.5$ and $f_{\text{mask, cat}}=0.2$. For (2), we train two separate models with all numerical features observed and $f_{\text{mask, cat}}$ set to 0.05 and 0.4, respectively. All Conditional TabDDPM models were trained on a single NVIDIA H100 GPU. The training time for the mixed-imputation model was 4.53 hours, and both the categorical-only variants took less than 3 hours.
@@ 4 Evaluation
Dataset: Our study utilizes the ResStock [49] dataset (https://resstock.nlr.gov/). ResStock is a synthetic dataset statistically representative of the entire residential building stock in the U.S [15]. This dataset is generated by sampling from probability distributions derived from multiple data sources [49], resulting in an extensive collection of U.S residential buildings at PUMA-level resolution [13, 11]. A PUMA varies in area, consists of many geographically close buildings, and contains granular details on building characteristics in the area. We use the 2024.1 release version of the ResStock dataset, which comprises 2.2 million buildings. Each row of this tabular data corresponds to a unique building, and the columns include various building characteristics and energy consumption details. From this list of available features, we select a small subset to work with:
@@
In total, we have 35 features (4 numerical and 31 categorical). We always keep the latitude and longitude columns known or observed in our conditional generation setup. We use a small subset of withheld PUMAs [15] to create an out-of-distribution (OOD) test dataset for evaluating our generative model’s performance in unseen geographical regions. Overall, we split the dataset into train (1,405,212), validation (351,303), test (439,128), and OOD-test set (3,665).
@@ 4.1.1 Univariate analysis
Result summary: The resulting true and generated distributions are visualized as histograms, and we show plots for both numerical and categorical features in Figure 2. Quantitative distances are reported in Table 1, showing an average over every unique dependency set for each variable. The Wasserstein distance is normalized by the range of the true samples. Table 1 includes results for the mixed-imputation model. We also compute results for one of the categorical-only imputation models, trained using a masking ratio of $0.05$ (which is equivalent to masking up to $2$ out of $31$ categorical features during training). The performance of the categorical-only imputation model remains consistent with that of the mixed imputation model, with an average JS-Distance of $0.197$ after averaging over all categorical variables. For comparison, we include a baseline that imputes each categorical variable by randomly choosing a category from all its possible categories. The JS distance for this baseline is averaged across all variables and reported in the last row of Table 1, which shows the random baseline’s poor performance as compared to our model. While deterministic predictive methods (such as gradient-boosted decision trees like XgBoost) are common tabular imputation baselines, they are unsuitable for our problem scope, as their fixed input-output mapping architectures would require training a computationally intractable number of distinct models to accommodate the arbitrary combinations of missing features present in a building characteristics dataset. Even the univariate and bivariate evaluations presented here on the ResStock dataset would require training hundreds of individual models, a complexity that scales into the millions as more features are simultaneously masked in real-world urban energy modeling setups (as seen in our case study).
@@
We observe from Table 1 and Figure 2 that our model demonstrates strong conditional generation quality across many features, though the performance varies, with some variables performing better than others. According to Table 1, some variables, such as heating/cooling Setpoint has offset, heating fuel, and building geometry type, are particularly challenging for the model to learn. This may be due to complex underlying class structures or imbalanced classes for some of these variables [11]. Figure 2 shows that the model’s output is realistic and matches the true distribution closely, even with features having a higher number of categories, e.g., for refrigerator, vintage (built year range), and water heater efficiency; the generated distribution also captures the dominant modes well, like "Natural Gas Standard" for water heater efficiency. Notably, the model performs well even on features that lack strongly defined dependencies in the ResStock data, such as the clothes dryer usage level variable. The model is also effective in capturing the wide range of values present in the numerical distribution, e.g, for square footage.
@@
Analysis on OOD: Here, we validate our model’s performance on the out-of-distribution (OOD) test dataset. Overall, we expect OOD conditional generation to be a challenging task, particularly for the variables that directly or indirectly depend on geospatial factors, as buildings in these PUMAs were not seen during training. The Figure 3 shows generated vs true distribution plots on a subset of the variables visualized in Figure 2. We see similar qualitative performance in Figure 3 for all these variables. It is important to note that for variables such as vintage or building square footage, which are dependent on PUMA (as seen in the dependencies obtained from ResStock), a comparison to the true distribution is not possible since there are no matching training samples with the same PUMA. However, this challenge highlights the advantage of our approach - learning a single generative model on the entire U.S. residential building stock. This enables our model to still generalize from broader patterns and impute building characteristics for geographical locations unseen during training, unlike studies that train a separate neural network model for each district ([13, 11]).
@@ 4.3 Case Study
We demonstrate the application of our framework in a case study on a residential neighborhood in Baltimore, Maryland (see Figure 5). Our study integrates our conditional generative modeling pipeline with the URBANopt [12, 36] energy model following the setup in Kontar et al. [11]. Evaluating our model’s performance at the community level is the most relevant metric for our intended application, as it allows us to capture energy use patterns at this scale and supports district-scale planning. Specifically, given a set of observed characteristics for the $77$ buildings in this neighborhood, we employ our Conditional TabDDPM model to impute all the unknown building characteristics. This synthetically completed data is then passed as an input to URBANopt, a physics-based energy model, which simulates load profiles in terms of energy or electricity consumption per building.
Figure 5: A residential community in Baltimore is used in our case study (based on [11]). There are $77$ single-family attached buildings. Building characteristics such as building type, attic type, foundation type, building area (square footage), built year, number of stories, and heating fuel are obtained for these buildings from sources like OpenStreetMap and Zillow/Redfin.
We utilize building characteristics for the buildings in this neighborhood as curated in El Kontar et al.’s work [11], and we refer to this as the reference dataset. These include building attributes collected from open-source databases and characteristics produced by their neural network-based predictive model (discussed in 2.1). We emphasize that this reference dataset of building characteristics does not contain any actual metered data, as that would be extremely challenging to obtain. It is a hybrid of critical characteristics such as building type, attic type, foundation type, building square footage, vintage, geometric stories, and heating fuel collected from real-world sources including OpenStreetMap and Zillow/Redfin, and the remaining characteristics predicted from the neural network model [11]. From this set of characteristics, we select the $35$ building features in the Conditional TabDDPM model setup. We keep all the numerical features as observed and mask varying levels of categorical features. We start by masking one categorical feature, then two features, and then progressively increase the number of masked features to $10$, to assess our model’s ability in imputing increasing levels and arbitrary combinations of unknown characteristics conditioned on the rest of the observed attributes. Crucially, we specifically included the masking of three real-world derived variables: vintage, geometric stories, and heating fuel (bolded in Figures 6 and 7), to ensure we evaluate the model’s ability in recovering such characteristics. The observed attributes used for conditioning notably always included the real-world derived building square footage, as numerical features were kept fully observed in this case study. After imputing the corresponding number of missing (masked) variables with our model, we evaluate the imputed features’ quality by comparing URBANopt’s resulting load profile using this generated dataset against the true load profile, which is simulated using the reference dataset from [11].
@@ 5 Discussion
The results presented in the paper successfully validate the use of a conditional diffusion framework in generating realistic building attributes and creating complete characteristics in building-level datasets. Our findings in Section 4 demonstrate the model’s ability to generate imputed features whose distributions closely match the true ResStock distribution for many building variables. However, a feature-level breakdown of the JS distance between the two distributions in Table 1 shows that a few variables are more challenging to learn than others. It is possible that some of these features, like building types and the number of occupants, which are critical conditioning variables for generating other variables [11], are not strongly determined by other observed features in return. Moreover, detailed analysis of some variables, including a few with the worst performance, such as heating/cooling setpoint has offset, heating fuel, and water heater fuel, reveals that our model learns an almost deterministic conditional mapping in these specific cases. The model consistently collapses to a single generated outcome when it is sampled conditioned on a given set of observed values, as seen in Figure 8. This behavior, is characterized as conditional distribution mode collapse. This means that, while the generated values remain highly sensitive to the input variables, and vary as those conditions change, for a particular input the model produces a Dirac delta probability distribution (all mass at a single output class). While these variables perform great at single-point prediction, scoring high in reconstruction accuracy, the probabilistic nature of the outputs is not captured well in these cases. Although we observe this behavior to be limited to only a few variables, it is possible that an improved diffusion model workflow could address this in future work.
@@
Overall, our evaluation focused on comparing the generated and true probability distributions in the univariate and bivariate conditional generation setting. While we don’t extend this analysis to higher-dimensional multivariate cases, we demonstrate our model’s multivariate conditional generation capabilities by simultaneously imputing up to $10$ building attributes, i.e, nearly a third of the total variables, in a real-world case study evaluation. Furthermore, our validation experiments with the OOD dataset show the model’s strong performance in unseen geographical regions, which is an important advantage of our global generative modeling framework. Finally, we showcase an end-to-end application where our model generates a completed building-level dataset needed as an input to the URBANopt energy model. Our case study on a small residential community analyzes the simulated energy load profiles for these buildings, which closely match a curated (reference) baseline. This study underscores our model’s potential in generating physically plausible and reliable datasets for practical use cases, showing its utility in enhancing building energy modeling workflows. However, it is important to note that the generated outputs represent statistically realistic samples from a conditional probability distribution, and not the exact true values of a specific physical building. Notably, our two-stage workflow: using generative modeling for filling in data gaps, followed by physics-based simulation offers additional advantages in terms of preserving physical interpretability and enabling the exploration of retrofit analyses and strategies. Future work can involve expanding case studies to more complex and larger communities.
@@ 6 Conclusion
Acknowledgements
This work was authored by the National Laboratory of the Rockies for the U.S. Department of Energy (DOE), operated under Contract No. DE-AC36-08GO28308. This work was supported by the Laboratory Directed Research and Development (LDRD) Program at the National Laboratory of the Rockies. The views expressed in the article do not necessarily represent the views of the DOE or the U.S. Government. The U.S. Government retains and the publisher, by accepting the article for publication, acknowledges that the U.S. Government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for U.S. Government purposes. This research was performed using computational resources sponsored by the U.S. Department of Energy’s Office of Critical Minerals and Energy Innovation and located at the National Laboratory of the Rockies.